Using the five step method please solve the following: Erin's age is 3 times warren's. In 4 years she will be twice as old as he will be. How old is each now.
Answer by ankor@dixie-net.com(22740)

Using the five step method please solve the following: Erin's age is 3 times warren's. In 4 years she will be twice as old as he will be. How old is each now?
:
I'm going to do it and hope we have 5 steps
:
Let e = Erins's present age
Let w = Warren's present age
;
Write an equation for each statement:
:
"Erin's age is 3 times warren's."
e = 3w
;
"In 4 years she will be twice as old as he will be."
e + 4 = 2(w+4)
e + 4 = 2w + 8
e = 2w + 8 - 4
e = 2w + 4
:
How old is each now?
substitute 3w for e in the above equation, find w
3w = 2w + 4
3w - 2w = 4
w = 4 yrs is Warren's present age
then
e = 3(4)
e = 12 yrs is Erin's present age
:
:
:
check solution in the statement:
"In 4 years she will be twice as old as he will be."
12 + 4 = 2(4+4)